OpenStudy (anonymous):

Use the properties of exponents to simplify (3^2+sqrt5)(3^2-sqrt5)

OpenStudy (anonymous):

(x+a)(x-a) = x^2 - a^2

OpenStudy (anonymous):

3^2 = 9 so 9^2 - 5 = 81-5 = 76

OpenStudy (anonymous):

do I do this just like I foil it

OpenStudy (anonymous):

You do*.

OpenStudy (anonymous):

*although foil is word I like to use. Distribute is a better word.

OpenStudy (anonymous):

ok, we just use that in our high school.